A is a reflexive angle where sin(A)=-5/13. Find cos(4A). Show your step-by-step calculation.
Answer by stanbon(75887)

A is a reflexive angle where sin(A)=-5/13. Find cos(4A). Show your step-by-step calculation.
Note:: 180 < A < 360
----
arcsin(-5/13) = -22.62 degrees = 337.38 degrees
-----
cos(4A) = cos(4(-22.62)) = cos(-90) = 0
-------------------------
Cheers,
Stan H.
-------------